### v1

@@ 2.3 Single-domain applications
Before conducting extensive multi-domain benchmarking, we first examine in-domain cases where both the material type and XC functional are represented in the training set. One such benchmark is Matbench Discovery [43], which reports F1, , and RMSD, each probing a different facet of reliability. The F1 score (harmonic mean of precision and recall) measures the accuracy of crystalline energy ranking; is the symmetric relative mean error (SRME) for lattice thermal conductivity (), reflecting the fidelity of curvature at the PES minimum; and RMSD quantifies structural deviations after relaxation relative to DFT. A combined performance score (CPS) summarizes these metrics. 7net-Omni.mpa achieves F1 = 0.889, = 0.265, and RMSD = 0.0639, yielding CPS = 0.849, which is slightly higher than 7net-ompa.mpa (0.845). Notably, values for 7net-Omni.omat24 and 7net-Omni.matpes are 0.253 and 0.243, respectively, lower than for 7net-Omni.mpa, which can be attributed to the inclusion of high-energy configurations in OMat24 and MatPES.
To extend beyond tests on pure crystals, we compute grain boundary (GB) energies for 327 configurations across 58 elemental metals [56], covering a range of misorientation angles as well as tilt and twist boundaries. We exclude 30 GBs with non-orthogonal planes relative to the -direction [15]. The GB energy () is computed following the definition in ref. [56], and performance is quantified as the SRME of . Figure S1a summarizes the performance of 7net-Omni compared with other leading uMLIPs. All models exhibit comparable performance, but 7net-Omni.mpa performs slightly better.
As another non-crystalline benchmark, we compute defect binding energies in steels. Previous DFT studies have examined interactions among carbon interstitials and vacancies, as well as interactions between transition-metal solute atoms [57, 58], which have been adopted as a benchmark for uMLIPs [59]. In this benchmark, we compute binding energies between carbon and a vacancy, as well as between transition-metal solutes (Ti, V, Cr, Mn, Co, Ni, Cu, Nb, and Mo). For all relaxations of defective structures, both atomic positions and lattice parameters are allowed to vary, while the cell shape is constrained to remain orthogonal [59]. Figures S1b,c compare the accuracy of uMLIPs for defect binding energies. Overall, all models achieve reasonable accuracy, with no single model showing a clear advantage.
@@
In Fig. S1d, we benchmark several uMLIPs, including 7net-Omni.spice, 7net-Omni.omol25, DPA.spice, and UMA.omol25 as multi-task models, alongside single-task models for organic molecules, such as MACE-OFF24-medium trained on the SPICE dataset [63, 28] (denoted MACE[spice]) and eSEN-sm-conserving trained on the OMol25 dataset [27] (denoted eSEN[omol]). Although the dispersion corrections differ between the functionals used in SPICE (B97M-D3) and OMol25 (B97M-V [64]), the results are compared on equal footing. As shown in Fig. S1d, all models achieve high accuracy for torsional barriers, with MAEs well below the chemical accuracy threshold of 1 kcal/mol (gray dashed line in Fig. S1d).
@@ 2.4 Cross-domain or cross-functional scenarios
The main motivation of this work is to develop a uMLIP that can simulate complex material systems in which distinct classes of materials interact. One example is semiconductor processing, such as ALD, where precursor molecules adsorb onto silicon-based substrates. In such applications, semilocal functionals, most commonly the PBE functional, are preferred to describe both molecular and solid-state systems simultaneously [41]. Within the present multi-task training framework, this requires effective knowledge transfer from PESs learned with hybrid functionals to those following the PBE functional for molecules. To examine this more closely, we first focus on molecular systems, either isolated or crystalline, computed at the PBE level, as shown in Figs. 2a-c.
@@
Next, we compute the reaction energy () for organometallic complexes. A total of 53 reactions involving 97 organometallic complexes were collected from the literature and reoptimized at the PBE-D3 level of theory [67, 68]. Figure 2b shows the MAE of over the benchmark set, indicating that 7net-Omni outperforms all other models. Both 7net-Omni and UMA include organometallic complexes in their training sets (OMol25), yet their cross-functional performances differ substantially. Interestingly, 7net-Omni.matpes is more accurate than 7net-Omni.mpa, despite the latter channel incorporating some organometallic complexes through DBS. Further analysis reveals that reactions with large errors in 7net-Omni.mpa predominantly involve Cr, Fe, Co, and Ni centers (Fig. S2), suggesting that this underperformance is related to the application of PBE+ in the MPtrj/sAlex database for partially filled 3 transition metals. A more detailed discussion is provided in Section 2.5.
To assess the description of intermolecular interactions, which are critical in organic semiconductors and molecular liquids, we investigate molecular crystals. A total of 86 molecular crystals are considered, comprising 23 from the X23 set [69, 70] and 67 from the BMCOS1 benchmark set [71], spanning systems from simple crystals to optoelectronic materials. Four crystals that overlap with the BMCOS1 set are excluded from the X23 subset to avoid redundancy. We compute the cohesive energy (), defined as the energy required to separate the crystal into its constituent molecules. The MAE of in Fig. 2c shows that 7net-Omni.mpa outperforms other models, including UMA.omc, which is explicitly trained on molecular crystals. Fig. S3 further compares errors in the equilibrium volumes of molecular crystals, showing trends consistent with those in Fig. 2c.
@@
Figure 2e summarizes the model accuracy, showing that 7net-Omni achieves substantial improvements over 7net-ompa. Among the single-task models, eSEN and ORB also perform competitively.
Metal-organic frameworks (MOFs), as hybrid organic–inorganic materials, offer vast structural tunability through the nearly limitless combinations of linkers and metal nodes [77]. However, their immense chemical space and large atomic numbers often make DFT-based screening impractical. Recent advances in uMLIPs provide a scalable alternative [78]. To assess the accuracy of 7net-Omni in MOF applications, we conducted four benchmark tasks covering both homogeneous properties (heat capacity) and heterogeneous cases (MOF-guest interactions).
Full data in Fig. S5 show that all models performed well in predicting heat capacities [79, 80], but accuracy in MOF-guest interactions varied widely depending on channels and training databases for some models (UMA, DPA, and eSEN). Therefore, throughout Fig. 2f, we display only accurate variants: 7net-Omni.mpa, UMA.omc, DPA.mp, 7net-ompa.mpa, eSEN[oam], ORB[omat], NequIP, GRACE[omat], and MACE[mpa]. Except for UMA.omc, all models include D3 corrections separately.
@@
In Fig. S5b, 7net-Omni.odac23 underperforms in without deformation in the repulsion region, even though this task is directly trained with the database related to MOF-gas interactions. We suspect that this behavior arises from the relatively loose DFT settings used in the ODAC23 dataset, which may have introduced noise [33]. In preliminary tests (not shown), replacing ODAC23 with ODAC25 significantly reduced the error.
In the lower-left region of Fig. 2f, we further evaluated adsorption with framework deformation using a DFT dataset from prior work [82], again focusing on single-molecule CO2 and H2O adsorption. In this scenario, 7net-Omni.mpa delivered the highest accuracy, surpassing the practical 0.1 eV MAE threshold for screening (gray dashed line in Fig. S5d) [82].
@@ 2.5 Metallic surfaces
We first investigate the adsorption of simple adsorbates (*H, *O, *OH, *CO) on noble-metal surfaces (Au, Ag, Cu, Pd, Pt). These systems are included in the OC20 dataset, computed with the RPBE functional, and are also sparsely sampled in the DBS (see Table S1). For each metal, we consider both the (100) and (111) surfaces, with three symmetric adsorption sites: bridge, hollow, and top for (100), and fcc, hcp, and top for (111). The adsorption energy of adsorbate X is computed using Eq. 6, where for H, O, and OH is determined from the equilibrium of H2 and H2O molecules, while for CO the molecular energy is used directly. In total, 120 adsorption energies were computed.
As shown in Fig. 4a, 7net-Omni significantly outperforms other models, achieving a low MAE of approximately 0.06 eV in . The white bullets on the multi-task models indicate the MAE of the oc20 channel relative to RPBE reference values, showing that the PBE performance of 7net-Omni exceeds that of 7net-Omni.oc20. This may be due to the much larger number of structures available at the PBE level compared with RPBE, which biases the model weights toward PBE fidelity. A similar trend is observed for DPA.omat. By contrast, UMA.omat shows much larger errors than UMA.oc20, indicating inefficient knowledge transfer. (Among various channels of UMA, only UMA.omat was able to produce reasonable results for tests in this section.) Single-task models, with the exception of eSEN[oam], fail to achieve comparable accuracy. This is primarily because of their limited ability to describe reference molecular energies, as discussed in the previous section, which leads to systematic shifts in the reaction energies of identical adsorbates. The large errors in eSEN[omat] in comparison with eSEN[oam] are attributed to the inaccurate energies for some molecules as shown in Fig. 3a.
To extend the evaluation to more diverse metals and adsorbates, we consider the ADS41 dataset [86, 87]. ADS41 comprises 15 physisorption systems (primarily organic molecules on noble-metal surfaces) and 26 chemisorption systems. Reference adsorption energies were computed at the PBE-D3 level of theory [87]. The benchmark results are shown in Figs. 4b-d. Figure 4b displays MAEs for physisorption energies, while Fig. 4c presents chemisorption energies excluding Co and Ni. The anomalous Co and Ni systems are displayed separately in Fig. 4d. Overall, Figs. 4b,c reveal consistent trends across uMLIPs, with 7net-Omni and eSEN[oam] performing best, in agreement with Fig. 4a.
In the previous section, we showed that destabilization of molecular energies significantly affected the PES. To examine such effect in catalytic reactions, we investigate the CO2RR process on Pt and Pd surfaces. Specifically, we consider COOH formation pathways: CO2 + *H → *COOH shown in Ref. [88] on Pt(111) and Pd(111) surface, where CO2 is initially physisorbed. Nudged elastic band (NEB) calculations are performed to identify the minimum-energy paths. The resulting PES profiles and error analysis in Fig. S8 show that the forward and backward barriers usually differ by 0.1–0.2 eV between DFT and uMLIPs except for 7net-Omni and eSEN[oam], indicating potential inaccuracies in the turn-over frequency of CO2RR reactions.
@@
The above discussion is relevant whenever transition metals and oxygen atoms are within the cutoff radius. For example, this occurs in CO binding to single-atom metal centers in organometallic complexes (Section 2.4) and in OH/CHO adsorption on metal nitrides (Fig. S4) [90]. These results underscore the need for caution when applying uMLIPs to adsorption on transition-metals.
@@
Recently, the meta-GGA regularized–restored Strongly Constrained and Appropriately Normed (r2SCAN) functional has attracted considerable attention, as it provides improved agreement with experimental results compared to conventional PBE calculations, particularly for properties such as thermal stability, phonons, and lattice volumes [91, 92]. However, r2SCAN computations are several times more expensive than with the PBE functional, and only recently large databases based on r2SCAN become available [52, 23, 31]. In training 7net-Omni, we included three databases at the r2SCAN level (see Table 1). For MP-r2SCAN, data were collected directly from the Materials Project database, while MatPES-r2SCAN and MP-ALOE are directly available in the public domain. The total number of r2SCAN entries (1,283,036) corresponds to only 0.8% of those at the PBE level.
@@
The benchmarking results for r2SCAN-based uMLIPs are presented in Table 2. We first compute energies of crystals whose SCAN values are provided in the Alexandria database [94]. The crystal structures are fixed and with respect to elemental phases is compared. For cross-domain applications involving surface and molecular configurations, we consider the ADS41 and BMCOS1 benchmark sets used in Section 2.4, which also provide r2SCAN data.
@@
As shown in Table 2, 7net-Omni exhibits the best overall accuracy except for the energy estimation of BMCOS1 dataset. The improved agreement with experimental for matpes-r2scan channel in comparison with mpa implies that the r2SCAN channel of 7net-Omni can be utilized to obtain thermal properties of materials more accurately than with the conventional PBE functional. Considering the high computational cost of r2SCAN calculations, this represents a significant advance in theoretical approaches.
@@
We also note in Table 2 that 7net-Omni.matpes-r2scan performs substantially better than MACE, despite both models being trained on largely the same r2SCAN databases. The additional MP-r2SCAN set for 7net-Omni contributes only near-equilibrium configurations with limited data size (see Table 1), so the performance gain can be attributed to effective transfer learning from the abundant PBE-level data with similar PES characteristics.
@@ 2.7 Inference speed
We compare the inference speed of various uMLIPs on an NVIDIA H100 GPU (94 GB) using MD simulations of carbon diamond structures. The performance results are summarized in Fig. S7, where the -axis denotes the number of atoms and the -axis represents simulation throughput in ns/day. For libraries accelerating tensor-product operations, 7net-Omni was tested with FlashTP [99], MACE with cuEquivariance (v0.6.1), and NequIP with OpenEquivariance (v0.4.1) [100]. For models supporting LAMMPS [101] execution, the measurements were performed using LAMMPS, while those without LAMMPS support (eSEN, UMA, ORB) were evaluated using ASE MD [102]. Although DPA supports LAMMPS, only ASE results are reported here due to installation difficulties. It is observed that 7net-Omni utilizing FlashTP achieves competitive throughput. By contrast, eSEN, which achieved accuracy comparable to 7net-Omni in many scenarios, requires a large amount of memory and exhibits very slow performance.

### v2

@@ 2.3 Single-domain applications
Before conducting extensive multi-domain benchmarking, we first examine in-domain cases where both the material type and XC functional are represented in the training set. One such benchmark is Matbench Discovery [42], which reports F1, , and RMSD, each probing a different facet of reliability. The F1 score (harmonic mean of precision and recall) measures the accuracy of crystalline energy ranking; is the symmetric relative mean error (SRME) for lattice thermal conductivity (), reflecting the fidelity of curvature at the PES minimum; and RMSD quantifies structural deviations after relaxation relative to DFT. A combined performance score (CPS) summarizes these metrics. 7net-Omni.mpa achieves F1 = 0.889, = 0.265, and RMSD = 0.0639, yielding CPS = 0.849, which is slightly higher than 7net-ompa.mpa (0.845). Notably, for 7net-Omni.omat24 and 7net-Omni.matpes are 0.253 and 0.243, respectively, lower than for 7net-Omni.mpa, which can be attributed to the inclusion of high-energy configurations in OMat24 and MatPES.
To extend beyond tests on pure crystals, we compute grain boundary (GB) energies for 327 configurations across 58 elemental metals [57], covering a range of misorientation angles as well as tilt and twist boundaries. We exclude 30 GBs with non-orthogonal planes relative to the -direction [15]. The GB energy () is computed following the definition in ref. [57], and performance is quantified as the SRME of . Supplementary Fig. 2a summarizes the performance of 7net-Omni compared with other leading uMLIPs. Parity plots for GB energy benchmark are also illustrated in Supplementary Fig. 3. All models exhibit comparable performance, but 7net-Omni.mpa performs slightly better.
As another non-crystalline benchmark, we compute defect binding energies in steels. Previous DFT studies have examined interactions among carbon interstitials and vacancies, as well as interactions between transition-metal solute atoms [58, 59], which have been adopted as a benchmark for uMLIPs [60]. In this benchmark, we compute binding energies between carbon and vacancy, as well as between transition-metal solutes (Ti, V, Cr, Mn, Co, Ni, Cu, Nb, and Mo). For all relaxations of defective structures, both atomic positions and lattice parameters are allowed to vary, while the cell shape is constrained to remain orthogonal. Supplementary Fig. 2b,c compare the accuracy of uMLIPs for defect binding energies, as well as their parity plots can be found in Supplementary Figs. 4,5. Overall, all models achieve reasonable accuracy, with no single model showing a clear advantage.
@@
In Supplementary Fig. 2d, we benchmark several uMLIPs, including 7net-Omni.spice, 7net-Omni.omol25, DPA.spice, and UMA.omol25 as multi-task models, alongside single-task models for organic molecules, such as MACE-OFF24-medium trained on the SPICE dataset [33, 27] (denoted MACE[spice]) and eSEN-sm-conserving trained on the OMol25 dataset [26] (denoted eSEN[omol]). (See Supplementary Fig. 6 for parity plots.) Although the dispersion corrections differ between the functionals used in SPICE (B97M-D3) and OMol25 (B97M-V [64]), the results are compared on equal footing. As shown in Supplementary Fig. 2d, all models achieve high accuracy for torsional barriers, with MAEs well below the chemical accuracy threshold of 1 kcal/mol (gray dashed line in Supplementary Fig. 2d).
@@ 2.4 Cross-domain or cross-functional scenarios
The main motivation of this work is to develop a uMLIP that can simulate complex material systems in which distinct classes of materials interact. One example is semiconductor processing, such as ALD, where precursor molecules adsorb onto silicon-based substrates. In such applications, semilocal functionals, most commonly the PBE functional, are preferred to describe both molecular and solid-state systems simultaneously [65, 66]. Within the present multi-task training framework, this requires effective knowledge transfer from PESs learned with hybrid functionals to those following the PBE functional for molecules. To examine this more closely, we first focus on molecular systems, either isolated or crystalline, computed at the PBE level, as shown in Figs. 2a–c. (For parity plots of each benchmark, see Supplementary Figs. 7–9.)
@@
In Supplementary Fig. 7, parity plots for torsion barriers are shown for all models. As the MAE increases, the data points become more scattered, indicating that the errors are more random rather than systematic. This increased randomness makes the predictions less reliable. This trend is consistently observed throughout this work: larger MAE values generally correspond to weaker correlations.
@@
Next, we compute the reaction energy () for organometallic complexes. A total of 53 reactions involving 97 organometallic complexes were collected from the literature and reoptimized at the PBE-D3 level of theory [69, 70]. Figure 2b shows the MAE of over the benchmark set, indicating that 7net-Omni outperforms all other models. Both 7net-Omni and UMA include organometallic complexes in their training sets (OMol25), yet their cross-functional performances differ substantially. Interestingly, 7net-Omni.matpes is more accurate than 7net-Omni.mpa, despite the latter channel incorporating some organometallic complexes through DBS. Further analysis reveals that reactions with large errors in 7net-Omni.mpa predominantly involve Cr, Fe, Co, and Ni centers (see Supplementary Fig. 8), suggesting that this underperformance is related to the application of PBE+ in the MPtrj/sAlex database for partially filled 3 transition metals. A more detailed discussion is provided in Section 2.5.
To assess the description of intermolecular interactions, which are critical in organic semiconductors and molecular liquids, we investigate molecular crystals. A total of 86 molecular crystals are considered, comprising 23 from the X23 set [71, 72] and 67 from the BMCOS1 benchmark set [73], spanning systems from simple crystals to optoelectronic materials. Four crystals that overlap with the BMCOS1 set are excluded from the X23 subset to avoid redundancy. We compute the cohesive energy (), defined as the energy required to separate the crystal into its constituent molecules. The MAE of in Fig. 2c shows that 7net-Omni.mpa outperforms other models, including UMA.omc, which is explicitly trained on molecular crystals. (For parity plots, see Supplementary Fig. 9.) Supplementary Figs. 10,11 further compare errors in the equilibrium volumes of molecular crystals, showing trends consistent with those in Fig. 2c.
@@
Figure 2e summarizes the model accuracy, showing that 7net-Omni achieves substantial improvements over 7net-ompa. Among the single-task models, eSEN and ORB also perform competitively. (For parity plots, see Supplementary Fig. 13.)
Metal-organic frameworks (MOFs), as hybrid organic-inorganic materials, offer vast structural tunability through the nearly limitless combinations of linkers and metal nodes [79]. However, their immense chemical space and large atomic numbers often make DFT-based screening impractical. Recent advances in uMLIPs provide a scalable alternative [80]. To assess the accuracy of 7net-Omni in MOF applications, we conducted four benchmark tasks covering both homogeneous properties (heat capacity) and heterogeneous cases (MOF-guest interactions).
Full data in Supplementary Fig. 14 show that all models performed well in predicting heat capacities [81, 82], producing MAEs less than 0.03 J/K/g, but accuracy in MOF-guest interactions varied widely depending on channels and training databases for some models (UMA, DPA, and eSEN). Therefore, throughout Fig. 2f, we display only accurate variants: 7net-Omni.mpa, UMA.omc, DPA.mp, 7net-ompa.mpa, eSEN[oam], ORB[omat], NequIP, GRACE[omat], and MACE[mpa]. Except for UMA.omc, all models include D3 corrections separately.
@@
In Supplementary Fig. 14b, 7net-Omni.odac23 underperforms in without deformation in the repulsion region, even though this task is directly trained with the database related to MOF-gas interactions. We suspect that this behavior arises from the relatively loose DFT settings used in the ODAC23 dataset, which may have introduced noise [32]. In preliminary tests (not shown), replacing ODAC23 with ODAC25 significantly reduced the error.
In the lower-left region of Fig. 2f, we further evaluated adsorption with framework deformation using a DFT dataset from prior work [84], again focusing on single-molecule CO2 and H2O adsorption. In this scenario, 7net-Omni.mpa delivered the highest accuracy, surpassing the practical 0.1 eV MAE threshold for screening (gray dashed line in Supplementary Fig. 14d) [84]. Parity plots for the MOF benchmarks are illustrated in Supplementary Figs. 15–19.
@@ 2.5 Metallic surfaces
We first investigate the adsorption of simple adsorbates (*H, *O, *OH, *CO) on noble-metal surfaces (Au, Ag, Cu, Pd, Pt). These systems are included in the OC20 dataset, computed with the RPBE functional, and are also sparsely sampled in the DBS (see Supplementary Table 1). For each metal, we consider both the (100) and (111) surfaces, with three symmetric adsorption sites: bridge, hollow, and top for (100) surfaces, and fcc, hcp, and top for (111) surfaces. The adsorption energy of adsorbate X is computed using Eq. 6, where for H, O, and OH is determined from the equilibrium of H2 and H2O molecules, while for CO the molecular energy is used directly. In total, 120 adsorption energies were computed.
As shown in Fig. 4a, 7net-Omni significantly outperforms other models, achieving a low MAE of approximately 0.06 eV in . (See Supplementary Fig. 21 for parity plots.) The white bullets on the multi-task models indicate the MAE of the oc20 channel relative to RPBE reference values, showing that the PBE performance of 7net-Omni exceeds that of 7net-Omni.oc20. This may be due to the much larger number of structures available at the PBE level compared with RPBE, which biases the model weights toward PBE fidelity. A similar trend is observed for DPA.omat. By contrast, UMA.omat shows much larger errors than UMA.oc20, indicating inefficient knowledge transfer. (Among various channels of UMA, only UMA.omat was able to produce reasonable results for tests in this section.) Single-task models, with the exception of eSEN[oam], fail to achieve comparable accuracy. This is primarily because of their limited ability to describe reference molecular energies, as discussed in the previous section, which leads to systematic shifts in the reaction energies of identical adsorbates. The large errors in eSEN[omat] in comparison with eSEN[oam] are attributed to the inaccurate energies for some molecules as shown in Fig. 3a.
[…]
In the previous section, we showed that destabilization of molecular energies significantly affected the PES. To examine such effect in catalytic reactions, we investigate the CO2RR process on Pt and Pd surfaces. Specifically, we consider COOH formation pathways: CO2 + *H → *COOH shown in ref. [90] on Pt(111) and Pd(111) surface, where CO2 is initially physisorbed. Nudged elastic band (NEB) calculations are performed to identify the minimum-energy paths. The resulting PES profiles and error analysis in Supplementary Fig. 23 show that the forward and backward reaction barriers usually differ by 0.1–0.2 eV between DFT and uMLIPs except for 7net-Omni and eSEN[oam], indicating potential inaccuracies in the turn-over frequency of CO2RR reactions.
@@
The above discussion is relevant whenever transition metals and oxygen atoms are within the cutoff radius. For example, this occurs in CO binding to single-atom metal centers in organometallic complexes (Section 2.4) and in OH/CHO adsorption on metal nitrides (Supplementary Fig. 24) [92]. These results underscore the need for caution when applying uMLIPs to adsorption on transition-metals.
@@
Recently, the meta-GGA regularized-restored Strongly Constrained and Appropriately Normed (r2SCAN) functional has attracted considerable attention, as it provides improved agreement with experimental results compared to conventional PBE calculations, particularly for properties such as thermal stability, phonons, and lattice volumes [93, 94]. However, r2SCAN computations are several times more expensive than with the PBE functional, and only recently large databases based on r2SCAN become available [52, 22, 30]. In training 7net-Omni, we included three databases at the r2SCAN level (see Table 1). For MP-r2SCAN, data were collected directly from the Materials Project database, while MatPES-r2SCAN and MP-ALOE are directly available in the public domain. The total number of r2SCAN entries (1,283,036) corresponds to only 0.8% of those at the PBE level.
@@
As shown in Table 2, 7net-Omni exhibits the best overall accuracy except for the energy estimation of BMCOS1 dataset. The improved agreement with experimental for matpes-r2scan channel in comparison with mpa implies that the r2SCAN channel of 7net-Omni can be utilized to obtain thermal properties of materials more accurately than with the conventional PBE functional. Considering the high computational cost of r2SCAN calculations, this represents a significant advance in theoretical approaches. Parity plots in benchmarking r2SCAN can be found in Supplementary Figs. 25–29.
@@
We also note in Table 2 that 7net-Omni.matpes-r2scan performs substantially better than MACE, despite both models being trained on largely the same r2SCAN databases. The additional MP-r2SCAN set for 7net-Omni contributes only near-equilibrium configurations with limited data size (see Table 1), so the performance gain can be attributed to effective transfer learning from the abundant PBE-level data with similar PES characteristics. This is also consistent with the data-efficient multi-fidelity framework [43].
@@ 2.7 Inference speed
We compare the inference speed of various uMLIPs on an NVIDIA H100 GPU (94 GB) using MD simulations of diamond Si structures. The performance results are summarized in Fig. 5, where the -axis denotes the number of atoms and the -axis represents the simulation throughput in nanoseconds per day when the time step is 1 femtosecond. For libraries accelerating tensor-product operations, 7net-Omni was tested with FlashTP [101] (Omni-flash) as well as cuEquivariance (v0.7.0) [102] (Omni-cueq), MACE with cuEquivariance (v0.7.0), and NequIP with OpenEquivariance (v0.4.1) [103]. For models supporting LAMMPS [104] execution, the measurements were performed within LAMMPS, whereas those without LAMMPS support (eSEN, UMA, ORB) were evaluated using ASE MD [105]. Although DPA supports LAMMPS, only the ASE results are reported here due to installation difficulties.
The original SevenNet interface integrated with LAMMPS was used for the FlashTP library [10], while coupling with the cuEquivariance library was achieved through the ML-IAP package within LAMMPS [106]. The lowest precision settings were adopted for ORB (float32-high) and UMA (tf32). For the UMA model, additional acceleration was achieved by applying torch.compile and pre-merging linear experts, while activation checkpointing was enabled for systems containing more than 1,000 atoms to fit the model into GPU memory [8]. All models were tested with increasing system sizes until the GPU memory limit was reached.
In Fig. 5, it is observed that 7net-Omni, when accelerated with cuEquivariance, achieves a competitive throughput of approximately 0.36 ns/day for a system containing 10,000 Si atoms, although this is roughly half the speed of MACE, the fastest model in the present benchmark. In contrast, eSEN, which attains accuracy comparable to that of 7net-Omni in many cases, requires significantly more memory and demonstrates substantially slower performance. While cuEquivariance provides superior inference performance for 7net-Omni compared to FlashTP acceleration in larger systems (3,000 atoms), its throughput decreases for smaller systems. This decline is likely due to GPU underutilization or additional auxiliary overhead, indicating potential for further optimization. Notably, GRACE maintains high efficiency even for small system sizes.